OpenStudy (calculusxy):

Evaluate the sqaure root of -25.

OpenStudy (calculusxy):

@agent0smith

OpenStudy (agent0smith):

\[\large \sqrt{-25} = \sqrt{-1*25} =\sqrt{-1 }\sqrt{25} \]

jhonyy9 (jhonyy9):

why so ? everybody know that sqrt(-1) = i so than will be isqrt25 = +/- 5i

OpenStudy (calculusxy):

@jhonyy9 It would be \(5i\) right?

jhonyy9 (jhonyy9):

+/- 5i because (+5i)^2 = -25 and (-5i)^2 = -25

OpenStudy (agent0smith):

It is 5i and only 5i.

OpenStudy (calculusxy):

Why can it be -5?

OpenStudy (calculusxy):

Oh okay. So it's only 5i

OpenStudy (agent0smith):

\[\large x^2 = -25\]is not the same thing as \[\large x = \sqrt{-25}\]

OpenStudy (calculusxy):

Thanks all for your help. I have a couple more that I need help with

OpenStudy (calculusxy):

Actually i have only one. \(-|-8|\) would equal to \(-8\) right?

TheSmartOne (thesmartone):

yup :)

TheSmartOne (thesmartone):

oops, I forgot I was viewing closed questions. >.<